OpenStudy (anonymous):

help me! quick stats question X f 5.5 4 15.5 10 25.5 18 35.5 25 45.5 30 55.4 3 i need to find Q1, Q2, Q3, D7 and P30

ganeshie8 (ganeshie8):

whats D7 ?

OpenStudy (anonymous):

deciles

OpenStudy (anonymous):

i have my exam tomorrow please help!

ganeshie8 (ganeshie8):

@amistre64

OpenStudy (amistre64):

deciles are divisions by 10 percentiles division by 100 quartiles division by 4 etc ... the process is the same in each case, just sub in the appropriate divider

OpenStudy (amistre64):

\[1~k-cile=\frac{\#~elements}{k}\] \[n^{th}~k-cile=n*\frac{\#~elements}{k}\] and round or average as needed

OpenStudy (amistre64):

for example:there are \(\sum f\) individual elements \[Q1=\frac{\sum f}{4})\]

OpenStudy (anonymous):

Q1= 90+1/4 = 22.75th term which means i have to take the 22 and 23 term?

OpenStudy (amistre64):

\[Q3=3*\frac{\sum f}{4}\]

OpenStudy (amistre64):

22.75 rnds up to 23 if we had gotten 22.0000 then we would average the 22 and 23 elements

OpenStudy (amistre64):

and thats Q1 = 90/4, rounded not 90+1/4 :)

OpenStudy (amistre64):

90/10 = 1D 9*7 = 63 so you would average the 63,64 elements ... but if they are the same value then thats a rather pointless step :)

OpenStudy (anonymous):

i have seen in many website the formula for Q as i* n+1/4

OpenStudy (amistre64):

Qk = k*n/4

OpenStudy (amistre64):

youd have to link the sites for me to see what you are trying to express :)

OpenStudy (anonymous):

ok wait

OpenStudy (anonymous):

http://www.alcula.com/calculators/statistics/quartiles/ here you go

OpenStudy (anonymous):

http://quartilesdecilespercentiles.blogspot.com/# also this one

OpenStudy (amistre64):

i see, they are starting their count at n=0 or n=1 and are adjusting for it

OpenStudy (anonymous):

now which formula should i follow?

OpenStudy (amistre64):

1 2 3 4 5 6 lets take this for example: the median is equal to Q2 is 3.5 Q2 = 2*6/4 = 3 ... the average of 3+4 is 3.5

OpenStudy (amistre64):

Q0 = 0*6/4 = 0 ... what is the 0th position? may be what they are trying to address

OpenStudy (anonymous):

i don't get it. like even from that formula i could get 3.5

OpenStudy (amistre64):

by their account: Q2 = the 2*7/4 ... out right so they seem to be combining the averaging step to a more concise form

OpenStudy (amistre64):

they have to adjust when it is not an integer ... but im just used to the way i was taught it

OpenStudy (amistre64):

the formula you use is the one your material gives you to work with. is the best advice i can give for that one

OpenStudy (amistre64):

this problem may not make it so complicated ... since the like the 14th and 15th elements are the same value.

OpenStudy (amistre64):

90/4 = Q1 = 22.xx, so the 23rd value elements 15 thru 32 are all the same value

OpenStudy (anonymous):

no it is not just about one question. i mean which one should use generally?

OpenStudy (amistre64):

Q2 = 2*90/4 = 45 5.5 4 4 15.5 10 14 25.5 18 32 35.5 25 57 .... 45 is in these 45.5 30 55.4 3

OpenStudy (amistre64):

generally i use the one i posted to start with

OpenStudy (amistre64):

D7 = 7*90/10 = 70 5.5 4 4 15.5 10 14 25.5 18 32 35.5 25 57 45.5 30 87 70 is in this set .... 55.4 3

OpenStudy (amistre64):

P30 = 30*90/100 = 27 which is in the third grouping regardless of interpolation

OpenStudy (anonymous):

yeah i got it. but i am still confused about quartile formula. it is used in so many other formulas as well that i am afraid my answer will go wrong

OpenStudy (amistre64):

forget the formula and look at what is going on in the process. you are dividing the data set into Quart=4 parts and picking the 1,2,3,4 Q

OpenStudy (amistre64):

how accurate you need to be is up to the person doing the grading and all sorts of interpolations can be generated to match the accuracy needed.

OpenStudy (anonymous):

so either formula could be used?

OpenStudy (amistre64):

yes, they are both expressing the same thing but differ on how accurate you want the outcome.

OpenStudy (amistre64):

in this specific instance, the interpolation doesnt matter since some number between 25.5 and 25.5 is going to be 25.5

OpenStudy (anonymous):

ok thank you so much and wish me luck for tomorrow's exam :)

OpenStudy (amistre64):

good luck ;)

OpenStudy (anonymous):

:)